find the slope intercept of the line satisfying the given conditions
perpendicular to x=-1/3y passing through p(-1,0)

Answer by stanbon(75887)

find the slope intercept of the line satisfying the given conditions
perpendicular to x=-1/3y passing through p(-1,0)
----
Find the slope of the given line:
x = (-1/3)y
solve for "y":
y = -3x
slope = -3
------
Any line perpendicular to that line must have slope = 1/3
-----
Form: y = mx + b
0 = (1/3)(-1) + b
b = 1/3
-----
Equation:
y = (1/3)x + (1/3)
---------------------
cheers,
stan H.
=============